# spacing=nonfrench Sy-FAR: Symmetry-based Fair Adversarial Robustness

### v1

@@ 3 Threat Model
We consider settings where adversaries produce adversarial examples to mislead ML-based image classifiers, particularly, ones based on neural networks. The adversaries we consider have white-box access to models, enabling them to efficiently produce worst-case perturbations using their knowledge of model weights, in contrast to black-box counterparts, which often slower and achieve lower attack success [12]. Moreover, we consider adversaries launching either untargeted or targeted attacks. Furthermore, we focus adversaries that use realistic attacks against face-recognition (specifically, using the eyeglass attack [45]). However, for better compatibility with prior work on fair adversarial robustness [18, 63], we also study adversaries that produce imperceptible adversarial perturbations bounded in -norm against object recognition. In both cases, the adversary’s goal is to maximize their success rate—i.e., achieving any misclassification in the case of untargeted attacks, or obtaining misclassification to target classes in the case of targeted attacks—thus maximally harming the integrity of the ML-based system.
On the other side, the defender aims to prevent adversaries’ attempts to mislead the ML model to the extent possible (i.e., maximizing robust accuracy) while also maximizing benign accuracy on clean inputs. Specifically, the defender induces adversarial robustness at training time, through a form of adversarial training. Crucially, in this work, we place strong emphasis on fair adversarial robustness. In particular, additionally to deterring attacks and correctly classifying benign samples, our defender also seeks to maximize fair source and target adversarial robustness, both at the individual class and group levels.
@@ 4 Sy-FAR’s Technical Approach
Our proposed symmetry-based regularization builds upon confusion-matrix analysis, which has been recognized as a foundation for fairness evaluation [16]. Here, however, we focus on confusion matrices over adversarial examples. We follow a standard adversarial training framework, where the objective balances benign accuracy and adversarial robustness, and extend it with a fairness regularizer. Unlike earlier fairness-oriented regularizers that rely on specialized optimization schemes [18, 51, 59], Sy-FAR is conceptually straightforward yet effective: it introduces a differentiable regularization term that explicitly enforces symmetry between class pairs, thereby counteracting asymmetry in adversarial settings. The intuition is that if examples of class can be perturbed to resemble class , then the reverse should hold as well—class similarity is fundamentally a symmetric relation. Framing fairness in this way provides a natural and tractable objective even when considering unbalanced or biased datasets. Moreover, the approach is computationally lightweight, as evaluating the regularizer requires only a single pass over the confusion matrix, making it well-suited for scalable adversarial training.
@@ 4.1 Intuition: Why Symmetry?
Fig. 2 illustrates two confusion matrices over adversarial examples. While one matrix is asymmetric and another is symmetric, both represent models with an overall 60% robust accuracy. Note, however, that the model with the asymmetric confusion matrix exhibits unfair source- and target-class adversarial robustness: class 2 has robust accuracy lower by 40% than class 1, and adversarial examples have 40% higher likelihood of being classified into class 1 than into class 2. In contrast, the model with the symmetric confusion matrix achieves fair source- and target-class adversarial robustness (both classes have identical robust accuracy and they are equally likely of being selected as output of successful adversarial examples). Thus, overall, this illustrative examples shows that if we are able to maintain (or increase) robust accuracy while improving symmetry, it is possible to promote fair adversarial robustness. In our experiments (§6–7), we find that symmetry regularization indeed preserves or boosts robust accuracy, and, as expected, also improved fair adversarial robustness from source- and target-class perspectives.
@@ 4.3 Training With Symmetry Regularization
where is calculates the cross-entropy loss on clean inputs and adversarial inputs ; is the soft confusion matrix of the adversarial batch; and are scalar hyperparameters balancing the three losses.
@@ 5 Experimental Setup
To evaluate fair adversarial robustness under realistic settings, we conduct experiments on three variations of the PubFig face-recognition dataset [23]. First, we use a gender-balanced subset of PubFig featuring ten celebrity identities (five female, five male), with approximately 300 images per identity.222Males: Antonio Banderas; Colin Powell; Hugh Grant; John Travolta; Will Smith. Females: Jennifer Lopez; Meryl Streep; Oprah Winfrey; Reese Witherspoon; Tyra Banks. This selection was motivated by the need to control for demographic imbalance: by ensuring equal male–female representation, differences in misclassification rates can be attributed to model behavior rather than dataset bias. As our main model, we use the VGG-16 architecture [49], a well-established convolutional neural network (CNN). Second, we modify the dataset to PubFig by replacing two female identities in PubFig with a pair of visually similar female siblings.333We replace Meryl Streep and Oprah Winfrey with Dakota and Elle Fanning, while retaining the remaining identities from PubFig. Doing so preserves the gender balance (five male, five female identities) and the dataset size (300 images per identity, as in PubFig), with the two replaced identities selected at random to maintain diversity. PubFig setup stress-tests the system’s ability to ensure fairness not only across gender but also among individuals with high visual similarity—a critical scenario for symmetry-aware fairness evaluation. Finally, we test generalization beyond CNNs by applying Sy-FAR to a Vision Transformer (ViT) trained on PubFig.
We pre-process all face images using the PyTorch-based FaceX-Zoo toolkit [56], which provides standard pipelines for face alignment and cropping. We adopt an 80/10/10 split for training, validation, and testing, a standard ratio that provides sufficient training diversity while keeping held-out data for reliable evaluation. Across all experiments, we perform 10 independent repetitions per setup and report averaged results, where each training method evaluated under its optimized hyperparameters. All methods are fine-tuned on top of adversarially trained models, where we empirically find that five epochs suffice for the PubFig and PubFig settings, while ten epochs yielded optimal results for PubFig.
@@
To enable direct comparison with prior work that focused primarily on CIFAR-10 and CIFAR-100 [22], and to demonstrate that our results generalize beyond face recognition, we also evaluate Sy-FAR on these benchmarks [22] with Preact-ResNet18 model, as commonly used in the adversarial robustness and fairness literature [59, 51, 63, 18]. Since these datasets do not capture the real-world fairness challenges central to our work (e.g., demographic balance or sibling similarity), we defer the detailed results in App. A. For both CIFAR-10 and CIFAR-100 with PreAct-ResNet18, all models are adversarially trained from scratch for 200 epochs.
@@
All face recognition experiments adopt the Rectangular Occlusion Attack (ROA) defense, following the FACESEC framework [54]. ROA simulates physical-world patch-based occlusions by performing an exhaustive search over fixed-size rectangular regions to find the most damaging location, then runs constrained attack within that region. We use FACESEC’s default parameters: rectangle size 7070, stride (10,10), step size of =20, and 100 attack iterations, with occlusion color initialized to mid-gray ( in pixel space). These settings give the best results in our experiments, offering strong robustness while maintaining competitive clean accuracy, and we adopt them consistently across all face-recognition setups.
On CIFAR-10 and CIFAR-100, we adversarially train the PreAct-ResNet18 model using PGD- [33], with perturbation budget =8/255, step size = 2/255, and ten iterations per step. This configuration is common in recent fair adversarial robustness studies, including FAAL [63] and SpecNorm [18], and thus provides a standard benchmark for comparison.
Fine-tuning for Enhancing Robust Fairness
Since our focus is specifically on the challenge of fair adversarial robustness, we build on adversarially trained models that already achieve a reasonable level of average robustness. This strategy is consistent with prior work, such as FAAL [63], which fine-tunes on top of adversarial training to further improve fairness-oriented objectives. In this context, we investigate whether fairness can be effectively improved by fine-tuning robust models, rather than retraining entirely from scratch. Our experiments indicate that adversarial fine-tuning is a more efficient approach: it enhances both fairness and robustness while avoiding the overhead of training from scratch.
@@
We evaluate the face-recognition models under the realistic eyeglass attack [45]. The attack overlays a fixed eyeglass-frame mask onto the face image and perturbs pixel values only within the masked region. The perturbation is initialized to a fixed color maximizing the (mis)classification loss. Adversarial examples are then refined through iterative gradient-based updates constrained to the eyeglass region, with momentum to stabilize optimization. We test a wide range of iterations capturing both weak and strong attack regimes; results are reported for the strongest attack configuratoin (wtih 300 iterations). Moreover, we use both the untargeted (primarily) and targeted variants of the attack. This setting reflects realistic threats where adversarial glasses can be realized by adversaries to mislead face recognition, making it a natural testbed for fairness evaluation. Fig. 3 depicts attack examples with images from PubFig, demonstrating how visually plausible perturbations can drastically impact recognition outcomes.
On CIFAR-10 and CIFAR-100, evaluation is performed using AutoAttack [12], a parameter-free ensemble of four strong adversarial attacks: (1) APGD-CE, (2) APGD-DLR [12], (3) FAB [11], and (4) Square Attack [2]. Consistent with prior fairness-aware adversarial training approaches, including FAAL[63] and SpecNorm[18], we adopt this evaluation setup as it provides a reliable benchmark for comparison.
@@
Source-class Fairness: We also evaluate fair adversarial robustness with respect to the source class. We adopt two standard metrics from the literature to quantify whether robustness is distributed evenly across classes [59, 51, 63, 18]: (1) Min (Worst-Class) Accuracy reports the lowest diagonal entry of the row-normalized confusion matrix, capturing the most vulnerable class under attack; and (2) Accuracy Gap measures the difference between the maximum and minimum values on the confusion matrix’s diagonal, measuring disparity across classes.
@@ 6 Class-Level Fairness: Experiment Results
In this section, we focus on untargeted eyeglass attacks against the face-recognition models. We note that the results on CIFAR-10 and CIFAR-100, reported in App. A follow consistent trends with those obtained on the face-recognition models. Similarly, results for the targeted eyeglass attack in App. B also corroborate our main findings.
@@ 6.1 Symmetry Loss
Diving deeper into the results shows that asymmetry in FAAL and SpecNorm stems from misclassifications between particular pairs of classes, where adversarial examples from one class are markedly more likely to be classified as the other class than vice versa. For instance, in the case of the PubFig dataset (Fig. 4), FAAL and SpecNorm are particularly asymmetric for classes 1 and 2 (Dakota and Elle Fanning, two sisters), 6 and 8 (Jennifer Lopez and Reese Witherspoon), and 6 and 9 (Jennifer Lopez and Beyoncé). By contrast, Sy-FAR yields higher symmetry, including for challenging classes where FAAL and SpecNorm struggle. Said differently, with Sy-FAR, pairwise confusions are more symmetric and no class disproportionately “loses” to another. Similar trends are observed for the CIFAR datasets (Table 5 in App. A) and targeted attacks (Table 8 and Fig. 8 in App. B).
@@ 6.2 Source-Class Robustness and Fairness
We next explore the effect of optimizing symmetry as part of Sy-FAR on fair source-class adversarial robustness as well as models’ overall accuracy and robustness. Specifically, we focus on benign accuracy, robust accuracy, worst-class robust accuracy, and robustness disparities between classes, as these metrics are the standard metrics in the literature on fair adversarial robustness [59, 51, 63, 18].
@@ 6.4 Run-time Comparison
We compare the computational cost of Sy-FAR against the baseline methods on the PubFig dataset using a VGG16 backbone. Experiments were conducted on a cluster node equipped with a single NVIDIA TITAN Xp GPU (12 GB), and the reported times are averaged across ten independent runs. Fig. 6 shows that Sy-FAR is substantially more efficient than FAAL and SpecNorm. By contrast, FAAL is nearly five times slower that Sy-FAR because of its bi-level optimization procedure that repeatedly solves inner adversarial updates for fairness-aware weighting. SpecNorm also incurs a significant slowdown—more than a factor of two—stemming from the need to compute the spectral norm of confusion matrix via singular value decomposition once per epoch. By comparison, standard adversarial training serves as the baseline cost (30 minutes/epoch), and Sy-FAR adds only a negligible overhead on top of it, since the operations from the confusion matrix (§4.3) are minor compared to the far larger cost of weight updates and adversarial example generation.
@@ 6.5 Stability Across Runs
Beyond average performance, a key requirement for defensive methods is stability: producing consistently strong results across different training runs. An unstable training method may require multiple attempts or fail to converge reliably under practical time and computational constraints. Indeed, we find that certain approaches are markedly more stable than others. We measure stability by examining the variance in robust accuracy, worst class accuracy and accuracy gap across ten independent runs. The results (shown in Figs. 9–11 in App. C) demonstrate that Sy-FAR achieves superior stability (i.e., lower variance) compared to alternative approaches not only in terms of robust accuracy, but also in terms of fair adversarial robustness. One possible explanation is that FAAL relies on bilevel optimization for fairness-aware weighting, which introduces additional non-convexity and sensitivity to initialization, while SpecNorm applies global spectral constraints that can interact unpredictably with adversarial updates. By comparison, Sy-FAR introduces a symmetry regularization term that, we posit, acts as a smoother objective, reducing run-to-run variance without adding significant optimization instability.

### v2

@@ 3 Threat Model
We consider settings where adversaries produce adversarial examples to mislead ML-based image classifiers, particularly, ones based on neural networks. The adversaries we consider have white-box access to models, enabling them to efficiently produce worst-case perturbations using their knowledge of model weights, in contrast to black-box counterparts, which often slower and achieve lower attack success [Croce20Auto]. Moreover, we consider adversaries launching either untargeted or targeted attacks. Furthermore, we focus on adversaries that use realistic attacks against face-recognition (specifically, using the eyeglass attack [sharif2016accessorize]). However, for better compatibility with prior work on fair adversarial robustness [jin2025enhancing, zhang2024towards], we also study adversaries that produce imperceptible adversarial perturbations bounded in -norm against object recognition. In both cases, the adversary’s goal is to maximize their success rate—i.e., achieving any misclassification in the case of untargeted attacks, or obtaining misclassification to target classes in the case of targeted attacks—thus maximally harming the integrity of the ML-based system.
Contrastively, the defender aims to prevent attempts to mislead the ML model to the extent possible (i.e., maximizing robust accuracy) while maximizing benign accuracy on clean inputs. Specifically, the defender induces adversarial robustness at training time, through a form of adversarial training. Crucially, in this work, we place strong emphasis on fair adversarial robustness. Specifically, additionally to deterring attacks and correctly classifying benign samples, our defender seeks to maximize fair source and target adversarial robustness, both at the individual class and group levels.
@@ 4 Sy-FAR’s Technical Approach
Our proposed symmetry-based regularization builds upon confusion-matrix analysis, which has been recognized as a foundation for fairness evaluation [gursoy2022equal]. Here, however, we focus on confusion matrices over adversarial examples. We follow a standard adversarial training framework, where the objective balances benign accuracy and adversarial robustness, and extend it with a fairness regularizer. Unlike earlier fairness-oriented regularizers that rely on specialized optimization schemes [jin2025enhancing, sun2023improving, xu2021robust], Sy-FAR is conceptually straightforward yet effective: it introduces a differentiable regularization term that explicitly enforces symmetry between class pairs, thereby counteracting asymmetry in adversarial settings. The intuition is that if examples of class can be perturbed to resemble class , then the reverse should hold as well—class similarity is fundamentally a symmetric relation. Framing fairness in this way provides a natural and tractable objective even when considering unbalanced or biased datasets. Moreover, the approach is computationally lightweight, as evaluating the regularizer requires only a single pass over the confusion matrix, making it well-suited for scalable adversarial training.
@@ 4.1 Intuition: Why Symmetry?
Fig. 2 illustrates two confusion matrices over adversarial examples. While one matrix is asymmetric and another is symmetric, both represent models with an overall 60% robust accuracy. Note, however, that the model with the asymmetric confusion matrix exhibits unfair source- and target-class adversarial robustness: class 2 has robust accuracy lower by 40% than class 1, and adversarial examples have 40% higher likelihood of being classified into class 1 than into class 2. In contrast, the model with the symmetric confusion matrix achieves fair source- and target-class adversarial robustness (both classes have identical robust accuracy and they are equally likely of being selected as output of successful adversarial examples). Thus, overall, this illustrative example shows that if we are able to maintain (or increase) robust accuracy while improving symmetry, it is possible to promote fair adversarial robustness. In our experiments (§6–7), we find that symmetry regularization indeed preserves or boosts robust accuracy, and, as expected, also improved fair adversarial robustness from source- and target-class perspectives.
@@ 4.3 Training With Symmetry Regularization
where is the cross-entropy loss on clean inputs and adversarial inputs ; is the soft confusion matrix of the adversarial batch; and are scalar hyperparameters balancing the three losses.
@@ 5 Experimental Setup
To evaluate fair adversarial robustness under realistic settings, we conduct experiments on three variations of the PubFig face-recognition dataset [attribute_classifiers]. First, we use a gender-balanced subset of PubFig featuring ten celebrity identities (five female, five male), with approximately 300 images per identity.111Males: Antonio Banderas; Colin Powell; Hugh Grant; John Travolta; Will Smith. Females: Jennifer Lopez; Meryl Streep; Oprah Winfrey; Reese Witherspoon; Tyra Banks. This selection was motivated by the need to control for demographic imbalance: by ensuring equal male–female representation, differences in misclassification rates can be attributed to model behavior rather than dataset bias. As our main model, we use the VGG-16 architecture [simonyan2015very], a well-established convolutional neural network (CNN). Second, we modify the dataset to PubFig by replacing two female identities in PubFig with a pair of visually similar female siblings.222We replace Meryl Streep and Oprah Winfrey with Dakota and Elle Fanning, while retaining the remaining identities from PubFig. Doing so preserves the gender balance (five male, five female identities) and the dataset size (300 images per identity, as in PubFig), with the two replaced identities selected at random to maintain diversity. PubFig setup stress-tests the system’s ability to ensure fairness not only across gender but also among individuals with high visual similarity—a critical scenario for symmetry-aware fairness evaluation. Finally, we test generalization beyond CNNs by applying Sy-FAR to a Vision Transformer (ViT) trained on PubFig.
We pre-process all face images using the PyTorch-based FaceX-Zoo toolkit [wang2021facex], which provides standard pipelines for face alignment and cropping. We adopt an 80/10/10 split for training, validation, and testing, a standard ratio that provides sufficient training diversity while keeping held-out data for reliable evaluation. Across all experiments, we perform ten independent repetitions per setup and report averaged results, where each training method is evaluated under its optimized hyperparameters (see App. E for details on the training method).
All methods are fine-tuned on top of adversarially trained models, where we empirically find that five epochs suffice for the PubFig and PubFig settings, while ten epochs yielded optimal results for PubFig.
@@
To enable direct comparison with prior work that focused primarily on CIFAR-10 and CIFAR-100 [krizhevsky2009learning], and to demonstrate that our results generalize beyond face recognition, we also evaluate Sy-FAR on these benchmarks [krizhevsky2009learning] with Preact-ResNet18 model, as commonly used in the adversarial robustness and fairness literature [xu2021robust, sun2023improving, zhang2024towards, jin2025enhancing]. Since these datasets do not capture the real-world fairness challenges central to our work (e.g., demographic balance or sibling similarity), we defer the detailed results in App. A. For both CIFAR-10 and CIFAR-100 with PreAct-ResNet18, all models are adversarially trained from scratch for 200 epochs.
@@
All face recognition experiments adopt the Rectangular Occlusion Attack (ROA) defense, following the FACESEC framework [tong2021facesec]. ROA simulates physical-world patch-based occlusions by performing an exhaustive search over fixed-size rectangular regions to find the most damaging location, then runs constrained attack within that region. We use FACESEC’s default parameters: rectangle size 7070, stride (10,10), step size of =20, and 100 attack iterations, with occlusion color initialized to mid-gray ( in pixel space). These settings give the best results in our experiments, offering strong robustness while maintaining competitive clean accuracy, and we adopt them consistently across all face-recognition setups.
On CIFAR-10 and CIFAR-100, we adversarially train the PreAct-ResNet18 model using PGD- [madry2018towards], with perturbation budget =8/255, step size = 2/255, and ten iterations per step. This configuration is common in recent fair adversarial robustness studies, including FAAL [zhang2024towards] and SpecNorm [jin2025enhancing], and thus provides a standard benchmark for comparison.
Fine-tuning for Enhancing Robust Fairness.
Since our focus is specifically on the challenge of fair adversarial robustness, we build on adversarially trained models that already achieve a reasonable level of average robustness. This strategy is consistent with prior work, such as FAAL [zhang2024towards], which fine-tunes on top of adversarial training to further improve fairness-oriented objectives. In this context, we investigate whether fairness can be effectively improved by fine-tuning robust models, rather than retraining entirely from scratch. Our experiments indicate that adversarial fine-tuning is a more efficient approach: it enhances both fairness and robustness while avoiding the overhead of training from scratch.
@@
We evaluate the face-recognition models under the realistic eyeglass attack [sharif2016accessorize]. The attack overlays a fixed eyeglass-frame mask onto the face image and perturbs pixel values only within the masked region. The perturbation is initialized to a fixed color maximizing the (mis)classification loss. Adversarial examples are then refined through iterative gradient-based updates constrained to the eyeglass region, with momentum to stabilize optimization. We test a wide range of iterations capturing both weak and strong attack regimes; results are reported for the strongest attack configuration (with 300 iterations). Moreover, we use both the untargeted (primarily) and targeted variants of the attack. This setting reflects realistic threats where adversarial glasses can be realized by adversaries to mislead face recognition, making it a natural testbed for fairness evaluation. Fig. 3 depicts attack examples with images from PubFig, demonstrating how visually plausible perturbations can drastically impact recognition outcomes.
On CIFAR-10 and CIFAR-100, evaluation is performed using AutoAttack [Croce20Auto], a parameter-free ensemble of four strong adversarial attacks: (1) APGD-CE, (2) APGD-DLR [Croce20Auto], (3) FAB [croce2020minimally], and (4) Square Attack [andriushchenko2020square]. Consistent with prior fairness-aware adversarial training approaches, including FAAL[zhang2024towards] and SpecNorm[jin2025enhancing], we adopt this evaluation setup as it provides a reliable benchmark for comparison.
@@
Robust Accuracy: The average accuracy under adversarial attack (untargeted or targeted, depending on the setup). This is the primary measure of adversarial robustness and is widely used in the literature (e.g., [madry2018towards, engstrom2018evaluating, kannan2018adversarial, zhang2020mixup, lee2020adversarial]).
Source-class Fairness: We also evaluate fair adversarial robustness with respect to the source class. We adopt two standard metrics from the literature to quantify whether robustness is distributed evenly across classes [xu2021robust, sun2023improving, zhang2024towards, jin2025enhancing]: (1) Min (Worst-Class) Accuracy reports the lowest diagonal entry of the row-normalized confusion matrix, capturing the most vulnerable class under attack; and (2) Accuracy Gap measures the difference between the maximum and minimum values on the confusion matrix’s diagonal, measuring disparity across classes.
@@ 6 Class-Level Fairness: Experiment Results
In this section, we focus on untargeted eyeglass attacks against the face-recognition models. We note that the results on CIFAR-10 and CIFAR-100, reported in App. A follow consistent trends with those obtained on the face-recognition models. Similarly, results for the targeted eyeglass attack (App. B) and powerful face-mask attacks [tong2021facesec] (App. G) also corroborate our main findings.
@@ 6.1 Symmetry Loss
A deeper analysis shows that asymmetry in FAAL and SpecNorm stems from misclassifications between particular class pairs, where adversarial examples from one class are markedly more likely to be classified as the other class than vice versa. For instance, in the case of the PubFig dataset (Fig. 4), FAAL and SpecNorm are particularly asymmetric for classes 1 and 2 (Dakota and Elle Fanning, two sisters), 6 and 8 (Jennifer Lopez and Reese Witherspoon), and 6 and 9 (Jennifer Lopez and Beyoncé). By contrast, Sy-FAR yields higher symmetry, including for challenging classes where FAAL and SpecNorm struggle. Said differently, with Sy-FAR, pairwise confusions are more symmetric and no class disproportionately “loses” to another. Similar trends are observed for the CIFAR datasets (Table 5 in App. A) and targeted attacks (Table 8 and Fig. 8 in App. B).
@@ 6.2 Source-Class Robustness and Fairness
We next explore the effect of optimizing symmetry as part of Sy-FAR on fair source-class adversarial robustness as well as models’ overall accuracy and robustness. Specifically, we focus on benign accuracy, robust accuracy, worst-class robust accuracy, and robustness disparities between classes, as these metrics are the standard metrics in the literature on fair adversarial robustness [xu2021robust, sun2023improving, zhang2024towards, jin2025enhancing].
@@ 6.4 Run-Time Comparison
We compare the computational cost of Sy-FAR against the baseline methods on the PubFig dataset using a VGG16 backbone. Experiments were conducted on a cluster node equipped with a single NVIDIA TITAN Xp GPU (12 GB), and the reported times are averaged across ten independent runs. Fig. 6 shows that Sy-FAR is substantially more efficient than FAAL and SpecNorm. By contrast, FAAL is nearly five times slower than Sy-FAR because of its bi-level optimization procedure that repeatedly solves inner adversarial updates for fairness-aware weighting. SpecNorm also incurs a significant slowdown—more than a factor of two—stemming from the need to compute the spectral norm of confusion matrix via singular value decomposition once per epoch. By comparison, standard adversarial training serves as the baseline cost (30 minutes/epoch), and Sy-FAR adds only a negligible overhead on top of it, since the operations from the confusion matrix (§4.3) are minor compared to the far larger cost of weight updates and adversarial example generation. Additional run-time analysis is included in App. F.
@@ 6.5 Stability Across Runs
Beyond average performance, a key requirement for defenses is stability: producing consistently strong results across different runs. An unstable training method may require multiple attempts or fail to converge reliably under practical time and computational constraints. Indeed, we find that certain approaches are markedly more stable than others. We measure stability by examining the variance in robust accuracy, worst class accuracy and accuracy gap across ten independent runs. The results (shown in Figs. 9–11 in App. C) demonstrate that Sy-FAR achieves superior stability (i.e., lower variance) compared to alternative approaches not only in terms of robust accuracy, but also in terms of fair adversarial robustness. One possible explanation is that FAAL relies on bilevel optimization for fairness-aware weighting, which introduces additional non-convexity and sensitivity to initialization, while SpecNorm applies global spectral constraints that can interact unpredictably with adversarial updates. By comparison, Sy-FAR introduces a symmetry regularization term that, we posit, acts as a smoother objective, reducing run-to-run variance without adding significant optimization instability.